The width of the book is .42 cm shorter than its length. The book's perimeter is 5.96 cm. What are the width and length in centimeters?
Answer by macston(5194)

L=Length; W=Width=L-.42 cm; P=perimeter=2(L+W)
P=2(L+W) divide each side by 2
P/2=L+W Substitute for W
P/2=L+L-.42 cm
P/2=2L-.42cm Add .42 cm to each side
(P/2)+.42 cm=2L Plug in values
(5.96 cm/2)+.42 cm=2L
2.98cm+.42cm=2L
3.4 cm=2L Divide each side by 2
1.7 cm=L ANSWER 1: Length is 1.7 cm
W=L-.42 cm=1.7cm -.42cm=1.28 cm ANSWER 2: Width is 1.28 cm
CHECK
P=2(L+W)
5.96cm=2(1.7cm+1.28cm)
5.96cm=2(2.98cm)
5.96cm=5.96cm